OpenStudy (anonymous):

Rewrite the rational exponent as a radical expression. 3 to the 2 over 3 power, to the 1 over 6 power

OpenStudy (anonymous):

I will fan and medal I promise

OpenStudy (janu16):

got hcies?

OpenStudy (janu16):

choices?

OpenStudy (janu16):

@saxprincess325 ?

OpenStudy (anonymous):

hold on I get them for you

OpenStudy (janu16):

ok np just wanted to make sure you were there

OpenStudy (janu16):

(32/3)1/6 (xA)B = xAB so we multiply the exponents 3(2/3)(1/6) 3(2/18) 31/9 x1/y = y√x 31/9 = 9√3 The answer is 9√3

OpenStudy (janu16):

i got 9 squareroot 3 so if you have that as a chocie than its correct

OpenStudy (anonymous):

thank you both

OpenStudy (janu16):

who both? lol

OpenStudy (janu16):

no problem!

OpenStudy (anonymous):

oops lol I meant thank you @Janu16 that always does that my computer keeps correcting me LOL

OpenStudy (anonymous):

I have one more can u help

OpenStudy (anonymous):

Explain how the Quotient of Powers was used to simplify this expression. 5 to the fourth power, over 25 = 52

OpenStudy (janu16):

yes! got options?

OpenStudy (anonymous):

By simplifying 25 to 52 to make both powers base five, and subtracting the exponents By simplifying 25 to 52 to make both powers base five, and adding the exponents By finding the quotient of the bases to be one fifth, and cancelling common factors By finding the quotient of the bases to be one fifth, and simplifying the expression

OpenStudy (janu16):

5^45^2=25 Because you have the same base 5 on numerator and denominator, you can subtract their power. 5^4−2=25 5^2=25 25=25 By simplifying 25 to 52 to make both powers base five, and subtracting the exponents

OpenStudy (janu16):

so A

OpenStudy (anonymous):

thank you so much

OpenStudy (janu16):

no rpoblem!

OpenStudy (anonymous):

you helped me on 3 square root 2^7 what was the answer again

OpenStudy (janu16):

um i dont remeber lol

OpenStudy (anonymous):

do you remember how to do this

OpenStudy (janu16):

what were the choices

OpenStudy (anonymous):

the cube root of 2 to the seventh power 2 to the 3 over 7 power 2 to the 7 over 3 power 221 24

OpenStudy (janu16):

i got (2^7) 1/3 in fraction form but you i think its b

OpenStudy (anonymous):

THANK YOU o you have helped so much@!!!!

OpenStudy (janu16):

if you take out the one it would be power 3

OpenStudy (janu16):

no rpoblem